# Atomic-Scale Mechanisms of Li-Ion Transport Mediated by Li10GeP2S12 in Composite Solid Polyethylene Oxide Electrolytes

### v1

@@ 3.2 Structural properties from MD simulation
We focused our investigations on determining the ability of Li ions on this surface to move along the Li–O bond chains formed. We propose a Li vacancy-driven mechanism, whereby the movement of the ions is influenced by adjacent vacancies. The barrier for the Li ion to move from its initial position to a newly formed vacancy was determined using the NEB method described in the computational details. We observed that the chemical environment on the surface of LGPS plays a key role in determining the minimum energy paths. The first structure investigated showed that it is feasible for Li to move from one oxygen atom to the next with a Li vacancy along the mPEO-TMS chain, provided that sufficient S atoms are present to form either a LiO–2S or LiO–3S bond. When atoms such as Ge are present along the path, the energy barrier is high since Li fails to form sufficient bonds with the LGPS surface.
@@
Our DFT calculations, therefore, indicate that Li-ion transport channels can exist at the mPEO-TMS|LGPS interface; however, they are not continuous. Li ions can move to adjacent vacancies if the chemical environment keeps the cation sufficiently bonded along the path. Thus, our DFT calculations indicate that Li-ion transport channels can have low barriers to Li-ion transport and contribute to ionic conductivity pathways.
4 Conclusion
This study elucidates how LGPS nanoparticle loading modulates Li-ion transport in composite solid polymer electrolytes. Molecular dynamics reveal a sharp increase in ionic conductivity up to 3.2 x% of LGPS, followed by a decrease governed by classical polymer-segmental and interface-driven transport mechanisms. The volcano-type curve at low LGPS wt% is in agreement with the experimental results. At higher loading levels (20% or more of LGPS), experiment shows a renewed increase in ionic conductivity that MD simulations fail to capture, suggesting a transition to quantum-enabled, vacancy-mediated Li hopping at the LGPS nanoparticle surface or through the bulk nanoparticles.
Optimal ionic transport in the LGPS-polymer electrolyte relies on engineered polymer-ceramic interfaces and effective nanoparticle dispersion, as excessive loading leads to agglomeration and hinders conductivity. DFT results demonstrate that specific chemical environments can lower the migration barrier, which can lead to more favorable transport channels consistent with findings from the MD simulations. The DFT calculations indicate that mPEO/LGPS migration barriers are dictated by the local sulfur and germanium environments. This offers new design pathways for the improved performance of composite polymer electrolytes.
See pages - of SI.pdf

### v2

@@ 2.2 Structural properties from MD simulation
We focused our investigations on determining the ability of Li ions on this surface to move along the Li–O bond chains formed. We propose a Li vacancy-driven mechanism, whereby the movement of the ions is influenced by adjacent vacancies. The surface Li migration mechanism investigated here assumes the presence of Li ion vacancies on the surface. We expect that the LGPS surface will have Li vacancies with little energy cost just as is present in LGPS bulk that is responsible for its high bulk ionic conductivity 28. In addition, the three Li released while grafting the mPEO-TMS units to the LGPS particles will also contribute to the presence of vacancies present on the surface. The barrier for the Li ion to move from its initial position to a newly formed vacancy was determined using the NEB method described in the computational details, and only a single mPEO-TMS chain was included in the Li migration calculations. We observed that the chemical environment on the surface of LGPS plays a key role in determining the minimum energy paths. The first structure investigated showed that it is feasible for Li to move from one oxygen atom to the next with a Li vacancy along the mPEO-TMS chain, provided that sufficient S atoms are present to form either a LiO–2S or LiO–3S bond. When atoms such as Ge are present along the path, the energy barrier is high since Li fails to form sufficient bonds with the LGPS surface.
@@
Our DFT calculations, therefore, indicate that Li-ion transport channels can exist at the mPEO-TMS|LGPS interface; however, they are not continuous. Li ions can move to adjacent vacancies if the chemical environment keeps the cation sufficiently bonded along the path. Thus, our DFT calculations indicate that Li-ion transport channels can have low barriers to Li-ion transport and contribute to ionic conductivity pathways. The low-barrier pathways we identified are primarily associated with sulfur-rich pathways, which are statistically abundant given that there are many more sulfur atoms in the LGPS crystal structure. In a realistic percolation network, these sites are not isolated but form quasi-continuous channels. Even if interrupted by occasional high-barrier (Ge-rich) sites, the percolating nature of the network should allow Li ions to detour through the abundant low-barrier S sites.
Conclusion
This study elucidates how LGPS nanoparticle loading modulates Li-ion transport in composite solid polymer electrolytes. Molecular dynamics reveal a sharp increase in ionic conductivity up to 3.2 x% of LGPS, followed by a decrease governed by classical polymer-segmental and interface-driven transport mechanisms. The volcano-type curve at low LGPS wt% is in agreement with the experimental results. At higher loading levels (20% or more of LGPS), experiment shows a renewed increase in ionic conductivity that MD simulations fail to capture, indicating a shift to a ceramic-dominated transport mechanism. DFT calculations elucidate that Li-ion migration proceeds via vacancy-driven hopping. This transport channel, distinct from the classical polymer dynamics, becomes dominant at high loadings where nanoparticles agglomerate to form percolated, bulk-like domains, driving the observed conductivity increase.
Optimal ionic transport in the LGPS-polymer electrolyte relies on engineered polymer-ceramic interfaces and effective nanoparticle dispersion, as excessive loading leads to agglomeration and hinders conductivity through the polymer electrolyte. DFT results demonstrate that specific chemical environments can lower the migration barrier, which can lead to more favorable transport channels consistent with findings from the MD simulations. The DFT calculations indicate that mPEO/LGPS migration barriers are dictated by the local sulfur and germanium environments. This offers new design pathways for the improved performance of composite polymer electrolytes.
Supporting Information
Details on the Molecular Dynamics simulation setup, including force fields and equilibration protocols; tables summarizing molecular masses and experimental electrolyte compositions; and additional figures displaying experimental ionic conductivity, radial distribution functions (RDFs) for PEO, Li-ion cluster analysis, and illustrations of interface modifications.
Acknowledgements
This research is supported by the Vehicle Technologies Office (VTO), Department of Energy (DOE), USA, through the Battery Materials Research (BMR) program. Argonne National Laboratory is operated for DOE by UChicago Argonne, LLC under the contract number DE-AC02-06CH11357. The authors gratefully acknowledge the computing resources provided on Bebop and Improv, high-performance computing clusters, operated by the Laboratory Computing Resource Center at Argonne National Laboratory.